# Harder or Different? Understanding Generalization of Audio Deepfake Detection

### v1

@@ 5.2 ASVspoof 2021 LA
5.3 ASVspoof 2021 DF and In-The-Wild
Table 3: Decomposition of performance gap between ASVspoof 2019 and ASVspoof 2021 DF. The performance gap can be attributed solely to ‘difference’ (red).
@@ 5.3 ASVspoof 2021 DF and In-The-Wild
Table 4: Decomposition of performance gap between ASVspoof 2019 and In-the-Wild. The performance gap can be attributed solely to ‘difference’ (red).
@@
This final experiment assesses the trade-off between hardness and difference new, unseen attacks. For this purpose, ASVspoof 2021 DF, containing partially unseen attacks, and In-the-Wild, encompassing completely unseen and more recent attacks, are utilized. Table 3 displays the results for ASVspoof 2021 DF, where an average performance gap of $14\%$ to $26\%$ in EER is observed, predominantly attributed to the ‘difference gap’. The evaluation on In-the-Wild, c.f. Table 4, reveals an even more substantial performance gap ($31\%$ to $78\%$), which is entirely due to the ‘difference gap’. Consequently, all models perceive the new attacks as a domain shift, characterized not by an escalation in difficulty or complexity but rather as encountering new, unseen, and distinct data and/or features.
@@ 6 Discussion and Implications
These results are astounding. First, we had expected data augmentation to increase ‘hardness’. However, for input length truncation and especially for compression and telephone encodings in ASVspoof 2021 LA, we observe a marked increase in ‘difference’, c.f. Section 5.2. To us, this is very surprising, since introducing more noise to the test data could be considered the archetype of increasing data difficulty. Second, whenever models are presented with new data, i.e. unseen attacks, the lack of generalization performance is attributed near-exclusively to ‘difference’. This is in stark contrast to similar experiments in the vision domain, where Lu et Al. evaluate the generalization of models trained on CIFAR-10 to the out-of-domain database CIFAR-10.2 [33]. Here, a ‘performance gap’ of about $15\%$ accuracy is, on average, decomposed into a ‘difference gap’ of $5\%$ and a ‘hardness gap’ of $10\%$. This more balanced decomposition is in stark contrast to our experiments on ASVspoof 2021 DF and In-the-Wild, where the ‘difference gap’ is the sole contributor to performance degradation.
@@
Addressing the challenge of ‘harder’, data could potentially be within reach by developing more sophisticated models [33]. However, ‘difference’ poses a more elusive challenge and may not be effectively addressed by merely increasing model complexity. Recall that machine learning fundamentally is pattern recognition. Patterns that fall outside the training data are, by definition, difficult for a model to comprehend, given that it derives all its knowledge from the statistical aggregation of input-output relationships present in the training data. Thus, the consistent prevalence of the ‘difference’ gap across all experiments, especially as the dominant factor in ASVspoof 2021 and In-the-Wild, might explain why achieving true out-of-domain generalization has been so problematic in previous research. It appears that deepfake-detection and anti-spoofing models are overly specific to their training data, with even minor deviations treated as outside the training data distribution (i.e. ‘different’).
@@
We hypothesize that models learn specific features that perform well for the current database, but which are so specific that they may not necessarily transfer well to other out-domain databases. An example from related work is the ‘silence-shortcut’ in the ASVspoof 2019 dataset [29]: Here, the models can use the length of leading silence to achieve near-perfect performance. The length of the silence correlates with the target variable in ASVspoof 2019 train; and even generalizes to the ASVspoof 2019 test data, but is obviously not a semantically meaningful feature that can be expected to work in the real world. It may be likely that there are other such ‘shortcuts’ [39] in voice anti-spoofing datasets, which may be harder for humans to identify, but which nevertheless are very predictive clues that are exploited by anti-spoofing models. The existence of such ‘shortcuts’ would align with our observation of the dominance of the ‘difference gap’, which is not nearly as pronounced in, for example, the vision domain [33].
@@
Therefore, while increasing model size by adding more parameters may prove beneficial in other domains, such as generating text with large language models, it might not be as effective in the realm of supervised classification for audio anti-spoofing and deepfake detection. At the same time, the use of self-supervised learning (SSL), exemplified by SSL-W2V2 and WhisperDF, appears to be a promising approach, as indicated by the overall smaller ‘performance’ gaps. This might be because such pre-trained models, when employed for voice anti-spoofing, usually are not fine-tuned. This could allow them to extract less specific, and thereby better generalizing features.
@@ 7 Future Work and Conclusion
This paper presents a study demonstrating that the performance gap in anti-spoofing generalization can be dissected into a ‘hardness’ and a ‘difference’ component. Our hypothesis is that current failures in generalization are predominantly due to the significant yet inadequately addressed influence of the ‘difference’ gap, suggesting that detection models overfit to the training distribution.
Future research should aim to deepen the understanding of this issue by systematically analyzing and breaking down errors across a wide range of audio augmentations (such as, potentially, augmentation by band-pass filters, compression, room impulse responses, Gaussian noise, time- and pitch-shifting, etc.), as well as between different attacks, e.g. between autoregressive and transformer-based TTS models. Such investigations could provide a more comprehensive view of the conditions under which models achieve generalization versus those in which they latch onto features highly specific to the training dataset. Such research may be more beneficial than current approaches, which simply increase model size, a strategy which may address only the ‘hardness’, but ignore the key ‘difference’ problem of voice authenticity classification.

### v3

@@ 5.2 Experiment 2: ASVspoof 2021 LA
5.3 Experiment 3, 4: ASVspoof 2021 DF & In-The-Wild
Table 3: Decomposition of performance gap between $D$ = ASVspoof 2019 and $D^{\prime}$ = ASVspoof 2021 DF. The performance gap can be attributed solely to ‘difference’ (red).
@@ 5.3 Experiment 3, 4: ASVspoof 2021 DF & In-The-Wild
Table 4: Decomposition of performance gap between $D$ = ASVspoof 2019 and $D^{\prime}$ = In-the-Wild. The performance gap can be attributed solely to ‘difference’ (red).
@@
The final experiment assesses the trade-off between hardness and difference for unseen attacks. For this purpose, testing is performed using the ASVspoof 2021 DF and the In-the-Wild databases. Table 3 displays results for the ASVspoof 2021 DF database, for which an average performance gap of $14\%$ to $26\%$ in EER is observed, and predominantly attributed to the ‘difference gap’. Results for the In-the-Wild database, c.f. Table 4, reveal an even more substantial performance gap ($31\%$ to $78\%$), which is entirely due to the ‘difference gap’. Consequently, all models perceive the new attacks as a domain shift, characterized not by an escalation in difficulty or complexity but rather as an exposure to previously unseen, distinct attacks with different characteristics.
@@ 6 Discussion and Implications
Results reported above are perhaps surprising. We expected both input length truncation and variability in compression and encoding to increase ‘hardness’ since the introduction of nuisance variability to the test data usually increases classification difficulty. Instead, results show a marked increase in ‘difference’.
@@
Second, whenever models are presented with new data, i.e. unseen attacks, the lack of generalization performance is also attributed near-exclusively to ‘difference’. This is in stark contrast to similar experiments in the vision domain, where Lu et al. evaluate the generalization of models trained on CIFAR-10 to the out-of-domain database CIFAR-10.2 [32]. Here, a ‘performance gap’ of about $15\%$ accuracy is, on average, decomposed into a ‘difference gap’ of $5\%$ and a ‘hardness gap’ of $10\%$. This more balanced decomposition is different from our experiments on ASVspoof 2021 DF and In-the-Wild, where the ‘difference gap’ is the sole contributor to performance degradation.
@@
Solving ‘harder’ databases might become feasible with the development of more advanced models [32]. However, ‘difference’ poses a more elusive challenge and may not be effectively addressed by merely increasing model capacity. Recall that machine learning is fundamentally pattern recognition. Patterns falling outside the training data are inherently difficult for a model to comprehend, as the model’s knowledge is derived entirely from the statistical aggregation of input-output relationships in the training data. Consequently, the consistent prevalence of the ‘difference’ gap across all experiments, particularly as the dominant factor for ASVspoof 2021 and In-the-Wild, may explain why achieving true out-of-domain generalization has been so problematic in previous research. It appears that deepfake detection and anti-spoofing models are overly specific to their training data, with even minor deviations being treated as outside the training data distribution (i.e., ‘different’).
@@
We induce that models learn specific features that perform well for the current database, but that are so specific that they don’t transfer well to other out-of-domain databases. An example from related work is the ‘silence-shortcut’ in the ASVspoof 2019 dataset [28]: Here, models can use the length of leading silence to achieve near-perfect performance. The length of the silence correlates with the class label; and even generalizes to the ASVspoof 2019 test data, but is obviously not a semantically meaningful feature that can be expected to work in the real world. It is likely that there are other such ‘shortcuts’ [38] in voice anti-spoofing datasets, which may be harder for humans to identify, but which nevertheless are very predictive clues that are exploited by anti-spoofing models. The existence of such ‘shortcuts’ would align with the dominance of the ‘difference gap’, which is not nearly as pronounced in, for example, the vision domain [32].
@@
Therefore, while increasing model capacity by adding more parameters may prove beneficial in other domains, such as generating text with large language models, it might not be as effective in the realm of supervised classification for audio anti-spoofing and deepfake detection. At the same time, the use of self-supervised learning (SSL), exemplified by SSL-W2V2 and WhisperDF, appears to be a promising approach, as indicated by the overall smaller ‘performance’ gaps. Possibly, such pre-trained models may extract less specific, and thereby better generalizing features.
@@ 7 Future Work and Conclusion
We present in this paper a study which demonstrates that the performance gap in anti-spoofing generalization can be decomposed into ‘hardness’ and ‘difference’ components. Our hypothesis is that poor generalization is due mainly to the substantial yet inadequately addressed influence of the ‘difference’ gap, indicating that detection models overly adapt to the training distribution.
Future research might systematically analyze and break down errors across a wide range of audio augmentations (such as, potentially, augmentation by band-pass filters, compression, room impulse responses, Gaussian noise, time- and pitch-shifting, etc.), as well as between different attacks, e.g. between autoregressive and transformer-based TTS models. Such investigations could provide a more comprehensive view of the conditions under which models generalize well instead of learning artefacts which are too specific to attacks observed in training data. Such research may be more beneficial than the current concentration upon increasing model capacity, a strategy which may address only the ‘hardness’, but ignore the key ‘difference’ problem in speech spoofing and deepfake detection.